# Designing a skilled soccer team for RoboCup: exploring Skill-Set-Primitives through reinforcement learning

### v1

@@ 9 Conclusion
This work introduces the skill set developed for FC Portugal, a team that has proven its capabilities by securing victory in the last two RoboCup 3D Soccer Simulation League World Championships. The methodology used to develop these skills is based on two concepts: a custom algorithm that extends PPO with symmetry-leveraging abilities and skill-set-primitives. These primitives are a subset of motion primitives, and capture common action patterns, making it easier to transition between different behaviors. Furthermore, they allow the policy to be simplified, improving sample efficiency and stability, while still retaining excellent performance in learning complex behaviors. Skill-set-primitives facilitate faster learning without introducing bias, apart from setting the initial exploration area within the solution space.
@@
Following that, we developed the locomotion set, a collection of skills based on one skill-set-primitive, which now forms the core of FC Portugal’s team. This set includes an omnidirectional walk that accelerates, decelerates, rotates, and changes direction seamlessly and automatically, while being controlled by a shallow neural network with one hidden layer containing 64 neurons. The Dribble skill allowed for unprecedented ball control and maneuverability, while the Push skill addressed collision-prone scenarios, increasing ball possession and game stability. We employed virtual agents to streamline the learning process in a multi-agent environment, enabling efficient learning within a reasonable timeframe.
Additionally, we have released the codebase of our team to the robotics community. This open-source approach allows other teams to build upon our work, fostering collaboration and innovation in the realm of robotic soccer.
@@ Acknowledgments
Appendix A Appendix
A.1 Parameters
@@ A.2 Sprint-Kick
step
1
Step countera [3]
0
1
count
1
Counter since last side switch
0
1
z
1
Head height prediction [3]
0
1
gyro
3
Gyroscope (x-forward, y-left, z-up)
(0,1,2)
(-1,1,-1)
dgyro
3
$\frac{d}{dt}\begin{bmatrix}\text{gyro}_{x}&\text{gyro}_{y}&\text{gyro}_{z}\end{bmatrix}$ for $\Delta t=0.02$ s
(0,1,2)
(-1,1,-1)
acc
3
Accelerometer (x-forward, y-left, z-up)
(0,1,2)
(1,-1,1)
dacc
3
$\frac{d}{dt}\begin{bmatrix}\text{acc}_{x}&\text{acc}_{y}&\text{acc}_{z}\end{bmatrix}$ for $\Delta t=0.02$ s
(0,1,2)
(1,-1,1)
feet
12
(l)eft and (r)ight feet sensors with (p)oint of origin and (f)orce vector: (lpx,lpy,lpz,lfx,lfy,lfz,rpx,rpy,rpz,rfx,rfy,rfz)
(6,7,8,9,10,11, 0,1,2,3,4,5)
(1,-$,1$,-1,1,1-1,1)
dfeet
12
$\frac{d}{dt}\begin{bmatrix}\text{feet}_{0}&\ldots&\text{feet}_{11}\end{bmatrix}$ for $\Delta t=0.02$ s
(6,7,8,9,10,11, 0,1,2,3,4,5)
(1,-$,1$,-1,1,1-1,1)
joints
20
All joint positions excluding head and toes
Sagittal reflectionb
Sagittal reflectionb
djoints
20
$\frac{d}{dt}\begin{bmatrix}\text{joints}_{0}&\ldots&\text{joints}_{19}\end{bmatrix}$ for $\Delta t=0.02$ s
Sagittal reflectionb
Sagittal reflectionb
targetc
1
Target direction relative to the robot’s torso
0
-1
balld
3
Ball position relative to the robot’s torso   (x-forward, y-left, z-up)
(0,1,2)
(1,-1,1)
a In Sprint Forward and Sprint & Rotate, the counter has an upper limit of 93. A limit prevents the neural network from encountering higher values during testing than during training. The specific value was optimized as a hyperparameter, as it influences the final behavior.
b The symmetry operation associated with this sagittal reflection depends on the joints’ order
c The target is only included in Sprint & Rotate
d The ball position is only included in Kick
Table 5: Sprint-Kick action space modification sequence: Sprint Forward, Sprint & Rotate, Kick
@@
1. Action
Raw target positions for 22 joints (includes toes, excludes head joints)
2. Symmetry
The positions are reflected with respect to the robot’s sagittal plane when the left leg is positioned as the front leg (see Section 6.1)
3. IIR Filter
An exponential moving average filter is applied with a smoothing factor of 0.5
4. Map
The raw values are mapped to the position ranges of each individual joint
5a. Bias
The initial joint positions are added as a permanent bias (only for Sprint Forward)
5b. Primitive
Add cycle capture primitive (only for Sprint & Rotate and Kick)
6. Conversion
First, predict current joint positions by combining the last action with previous positions (the server has a 1-step delay). Then, determine joint speeds through proportional control.

### v2

@@ 7 Results and Discussion
8 Locomotion Performance in RoboCup
In the previous section, we presented results for individual skills tested in controlled conditions. While kicks are relatively straightforward to compare with published data and are often used similarly in training and competition, locomotion is heavily influenced by team strategies and must adapt to unforeseen circumstances during competitions. Moreover, comparing locomotion with published data is challenging due to the high variance in target versatility and performance evaluation metrics. To address these challenges, this section presents a statistical analysis of locomotion performance in past RoboCup competitions.
8.1 Methodology
As highlighted in Section 1, the NAO robot has been the standard platform in RoboCup Soccer since 2008. Consequently, our analysis begins with data from 2008, encompassing all major RoboCup competitions through 2023. Match data were sourced from the RoboCup.info online archive [18], and we included all rounds of each competition, including seeding rounds. Penalty shootouts, FatProxy matches, challenges, and test games were excluded. For each match, only data from the PlayOn segments was analyzed, as these periods impose fewer restrictions on player movement.
To ensure consistent and reliable data, we analyzed robot movement using 1-second measurement windows, which mitigates the impact of erratic movements or collisions. Specific exclusion criteria were applied to ensure the validity of these samples:
  • Abnormal height: Samples were excluded if the robot’s z-axis height was below 0.25 m or above 0.40 m.
  • Excessive speed: Samples were excluded if the robot’s speed between consecutive time steps exceeded 4 m/s, a threshold slightly higher than the maximum running speed.
Exclusion zones were extended 1 second prior to an invalid time step to account for events like falling, which could affect speed calculations. For example, if a robot begins to fall but is beamed by the referee before its height drops below 0.25 m — due to a play mode change or foul — the fall itself may go undetected, yet the initial falling motion would still be incorrectly factored into speed calculations.
8.2 Results
Table 3 summarizes the average and maximum speeds of RoboCup teams since 2008, aggregated by competition. Teams active since 2021 are ordered by the number of participations, with winners highlighted in red [^8] For completeness, SEU-RedSun won in both 2008 and 2009. In 2008, their average and maximum speeds were 0.12 m/s and 0.61 m/s, with a 15% fallen time and 41 m of average distance per player. In 2009, these values improved to 0.16 m/s, 0.73 m/s, 15%, and 56 m, respectively.. The key metrics were obtained as follows:
  • Average speed: computed by aggregating valid 1-second window samples for all teammates in a match, summing the distances covered, and dividing by the total window durations. This method is preferred over averaging individual speeds, as it gives more weight to players with more valid samples, preventing those who frequently fell or crashed from skewing the results. The process is repeated across all games and averaged per team.
  • Maximum speed: determined by selecting the 5th highest speed from all valid samples across all teammates in a match, ensuring the value reflects sustained performance rather than statistical anomalies.
From preliminary analysis, team speed appears positively correlated with competition success. This observation will be explored in greater detail later in this section.
Table 2: Average speed (m/s) and maximum speed (m/s) aggregated by team since 2008. Teams active since 2021 are ordered by number of participations; winners highlighted in red.
    FC Portugal     magmaOffenburg     UT Austin Villa     BahiaRT     HfutEngine3D     ITAndroids     Apollo3D     Miracle3D     KgpKubs     WITS FC     WrightOcean     MIRG
2008 .03  0.28 - .02  0.17 - .07  0.44 - .11  0.81 - - - - -
2009 .09  0.76 .05  0.26 - .02  0.10 .15  0.65 - - - - - - -
2010 .19  0.84 .14  0.46 .12  0.75 .03  0.37 .18  0.60 - .23  0.82 - - - - -
2011 .25  0.83 .16  0.67 .26  0.74 .05  0.54 .17  0.73 - .28  0.86 - - - - -
2012 .25  0.84 .20  0.73 .25  0.77 - - - .33  0.93 - - - - -
2013 .35  0.73 .28  0.87 .30  0.95 .11  0.45 .20  0.72 .19  0.71 .37  0.94 - - - - -
2014 .38  0.78 .35  0.86 .35  0.97 .31  0.86 .17  0.92 - - - - - - -
2015 .38  0.85 .35  0.85 .38  0.97 .36  0.77 .16  0.92 .26  0.63 .38  0.95 .33  0.61 - - - -
2016 .39  0.87 .37  0.85 .44  0.97 .34  0.74 .20  0.86 .27  0.76 - .35  0.61 .28  0.67 - - -
2017 .32  0.88 .39  0.86 .45  0.98 .39  0.87 .36  0.92 .31  0.70 - .20  0.84 .23  0.68 - - -
2018 .36  0.88 .38  1.05 .41  0.97 .39  0.85 - .31  0.70 - .21  0.78 .24  0.67 - - -
2019 .34  2.62 .42  1.06 .44  0.97 .41  0.84 .33  0.83 .34  0.70 - - - - .43  0.89 -
2021 .38  3.44 .44  1.07 .47  0.98 .37  0.85 .41  0.91 .33  0.70 .42  1.27 .41  0.79 .26  0.73 .27  0.78 .47  0.90 .27  0.78
2022 .48  1.43 .39  2.46 .47  0.98 .37  0.95 .40  0.92 .33  0.71 .43  0.82 .40  0.79 .32  0.84 .29  0.78 - -
2023 .46  0.93 .39  2.55 .31  1.64 .38  0.88 - .34  0.70 - - - .28  0.83 - -
Table 3: Fallen time (%) and average distance per player (m) aggregated by team since 2008. Teams active since 2021 are ordered by number of participations; winners highlighted in red.
    FC Portugal     magmaOffenburg     UT Austin Villa     BahiaRT     HfutEngine3D     ITAndroids     Apollo3D     Miracle3D     KgpKubs     WITS FC     WrightOcean     MIRG
2008 10% 130 - 22% 800 - 27% 250 - 14% 410 - - - - -
2009 32% 270 06% 210 - 27% 700 13% 520 - - - - - - -
2010 09% 880 12% 630 06% 650 33% 110 08% 860 - 11% 111 - - - - -
2011 11% 113 05% 800 05% 122 27% 150 08% 760 - 08% 128 - - - - -
2012 10% 111 03% 101 10% 114 - - - 21% 124 - - - - -
2013 08% 159 06% 136 04% 143 09% 480 05% 960 14% 750 07% 171 - - - - -
2014 08% 193 05% 155 06% 144 07% 169 08% 670 - - - - - - -
2015 06% 169 04% 168 03% 172 04% 163 09% 650 13% 100 09% 185 19% 118 - - - -
2016 05% 174 03% 180 03% 193 04% 168 09% 830 06% 123 - 10% 155 04% 139 - - -
2017 05% 156 03% 199 03% 194 06% 185 14% 129 06% 154 - 04% 970 04% 119 - - -
2018 06% 168 04% 185 03% 186 06% 181 - 07% 129 - 04% 990 04% 120 - - -
2019 10% 125 04% 190 03% 184 07% 177 05% 147 05% 140 - - - - 04% 161 -
2021 11% 134 03% 187 02% 195 06% 160 03% 159 04% 139 03% 180 04% 187 06% 120 03% 111 04% 166 04% 111
2022 03% 216 04% 164 03% 207 07% 150 03% 172 03% 133 02% 186 03% 179 06% 137 07% 990 - -
2023 03% 205 03% 139 06% 860 06% 148 - 04% 140 - - - 06% 102 - -
Table 3 reports the percentage of time robots were fallen and the average distance covered per player, aggregated by team since 2008. The fallen time metric reflects the proportion of PlayOn time during which a robot’s height was below 0.25 m. Average distances were calculated similarly to average speeds but normalized by the number of players per team, accounting for variations in team sizes over the years.
Note that total distance and average speed are not redundant metrics because the proportion of valid samples per match varies according to different factors. Consequently, a team may achieve a high average speed while covering a low total distance if its players fall frequently or matches are interrupted often.
Since the soccer field expansion in 2012, FC Portugal had prioritized locomotion stability over speed. In 2019, we introduced the league’s first running behaviors [3], including a sprint at 2.62 m/s with a slow turning rate of 10°10°10  °/s and a more agile run peaking at 1.52 m/s and 160°160°160  °/s [4]. However, faster behaviors proved harder to control, as the team’s strategy and path planning were not adapted to the drastic changes, leading to increased collisions, fallen time, and overall instability. In 2021, the team introduced the Sprint-Kick (Section 5.1), stabilizing at 3.44 m/s with a maximum turning rate of 33°33°33  °/s. The fallen time did not increase beyond 11% because the skill was only used during kickoffs.
Curiously, teams that introduced running behaviors experienced a reduction in the average speed. For FC Portugal (2019) and UT Austin Villa (2023), this was due to instability. In contrast, magmaOffenburg (2022) maintained stability, suggesting their reduced speed and distance were due to deliberate strategy changes.
In 2022, we introduced our new codebase with a dribbling skill peaking at 1.43 m/s. The fallen time was reduced from 11% to 3%, despite a significant increase in covered distance. Maximum walking speed improved from 0.88 m/s to 0.93 m/s, combined with unmatched omnidirectional capabilities (see Section 8.2.2). Then, in 2023, the fast dribbling skill was removed, as explained in Section 1.
8.2.1 Performance Correlation
To analyze the significance of the previous metrics on team performance, we calculated the Spearman rank correlation coefficients between these metrics and team rankings across all competitions (see Table 4). Team rankings, provided in Appendix A, indicate relative positions without reflecting the magnitude of performance differences. Unlike Pearson’s correlation, which assumes linear relationships, we used Spearman’s rank correlation to assess the strength and direction of the monotonic relationship between variables based on their ranks. Note that in competitions before 2015, some teams share the same rank. For example, in 2008, there are 7 rank groups, with the last five teams (ranks 17-21) all assigned rank 7 for the correlation calculation.
The average correlation coefficients across all competitions were 0.81 for average speed, 0.60 for maximum speed, 0.81 for covered distance, and --0.49 for fallen time. As expected, the maximum speed is a weaker predictor of performance compared to average speed or covered distance. Notably, in 2019 and 2021, FC Portugal’s maximum speed was significantly higher than other teams, yet it ranked 6th in both competitions.
The fallen time metric shows a negative correlation, indicating that greater fallen times, which suggest lower stability, are associated with lower rankings. While the magnitude of this correlation is weaker compared to other metrics, it still points to a meaningful relationship. The 2012 competition, however, was an outlier, where this trend was inverted, with teams exhibiting higher fallen times ranking better.
Table 4: Spearman rank correlation coefficients between team rankings and their performance metrics across years
2008 2009 2010 2011 2012 2013 2014 2015 2016 2017 2018 2019 2021 2022 2023
Avg. Sp. 0.81 0.72 0.86 0.86 0.70 0.90 0.79 0.79 0.87 0.75 0.79 0.82 0.92 0.79 0.77
Max. Sp. 0.74 0.58 0.69 0.60 0.38 0.62 0.74 0.52 0.50 0.39 0.86 0.43 0.76 0.61 0.66
Distance 0.80 0.89 0.88 0.89 0.88 0.89 0.65 0.80 0.88 0.85 0.86 0.89 0.85 0.78 0.31
Fallen T. -0.27 -0.32 -0.41 -0.46 0.21 -0.29 -0.62 -0.77 -0.50 -0.53 -0.61 -0.82 -0.54 -0.62 -0.83
8.2.2 Directional Analysis
In RoboCup competitions, robots mostly walk forward; however, walking backward or sideways is sometimes required, such as when aligning for a shot. To better characterize how teams employ omnidirectional locomotion, we examined how players move in 8 directions: forward, backward, sideways left and right, and the intermediate directions.
Measurements were taken by analyzing 1-second windows where speeds exceeded 0.1 m/s — below this threshold, intentional direction is hard to determine. Displacement vectors were compared to the robot’s average orientation during each window to assign one of 8 movement directions.
Fig. 9 shows directional metrics for FC Portugal’s locomotion during periods of significant evolution (2008–2013, 2017–2023): average speed (m/s), maximum speed (m/s), and distance covered (hm). The radial plots represent a top view of the robot, with the robot’s forward direction pointing upward. Data points are positioned radially to represent the direction of movement: forward at the top, backward at the bottom, leftward at the left, and so on.
Figure 9: FC Portugal locomotion statistics for periods of notable change (2008–2013, 2017–2023). Metrics include average speed (m/s), maximum speed (m/s), and distance covered (hm), plotted radially to show variation by movement direction, relative to the robot. Values at the top represent forward movement, values on the left indicate sideways movement to the left, and so on.
Initially, the team focused on developing a forward-walking skill, gradually introducing omnidirectional capabilities. Faster forward-running skills were introduced in 2019 and 2021. Asymmetries in the maximum speed arise from biases in these skills, particularly when stopping from a running state [3]. In 2022, the maximum forward speed was influenced by the dribbling skill. However, in 2023 that metric was solely dictated by the omnidirectional walk, revealing its consistent performance in all directions. Moreover, the distance metric indicates that omnidirectional walking was scarcely used until 2022, with back-walking between 2010 and 2012 being an exception.
Fig. 10 illustrates the same metrics for the six teams participating in the 2023 competition, arranged in order of their rankings; and Fig. 11 aggregates speed metrics for all teams over the years, including average speed, average maximum speed, and maximum speed (all in m/s).
Figure 10: 3DSSL locomotion statistics for the six teams that participated in RoboCup 2023, ordered by competition ranking. Metrics include average speed (m/s), maximum speed (m/s), and distance covered (hm), plotted radially to show variation by movement direction, relative to the robot. Values at the top represent forward movement, values on the left indicate sideways movement to the left, and so on.
Figure 11: 3DSSL aggregated locomotion statistics for periods of notable change (2008–2013, 2017–2023). Metrics include average speed, average maximum speed, and maximum speed (all in m/s), plotted radially to show variation by movement direction, relative to the robot. Values at the top represent forward movement, values on the left indicate sideways movement to the left, and so on.
In 2023, the two fastest teams in forward speed were magmaOffenburg and UT Austin Villa. Their skill asymmetries (Fig. 10) are apparent in replays [18]. Interestingly, these asymmetries are in opposite directions, complementing each other to create a symmetrical pattern in the aggregate plot (Fig. 11).
Although the maximum speed metric in the 3DSSL shows minimal apparent improvement from 2008 to 2018, robots fell significantly more often during earlier competitions, limiting the sustainability of their top speed and, consequently, their effectiveness. This conclusion becomes apparent when comparing the average speed metric over the same period.
@@ 10 Conclusion
This work introduces the skill set developed for FC Portugal, a team that has proven its capabilities by winning the last two RoboCup 3D Soccer Simulation League World Championships. The methodology used to develop these skills is based on two concepts: a custom algorithm that extends PPO with symmetry-leveraging abilities and skill-set-primitives. These primitives are a subset of motion primitives, and capture common action patterns, making it easier to transition between different behaviors. Furthermore, they allow the policy to be simplified, improving sample efficiency and stability, while still retaining excellent performance in learning complex behaviors. Skill-set-primitives facilitate faster learning without introducing bias, apart from setting the initial exploration area within the solution space.
@@
Following that, we developed the locomotion set, a collection of skills based on one skill-set-primitive, which now forms the core of FC Portugal’s team. This set includes an omnidirectional walk that accelerates, decelerates, rotates, and changes direction seamlessly and automatically, while being controlled by a shallow neural network with a single hidden layer of 64 neurons. The Dribble skill allowed for unprecedented ball control and maneuverability, while the Push skill addressed collision-prone scenarios, increasing ball possession and game stability. We employed virtual agents to streamline the learning process in a multi-agent environment, enabling efficient learning within a reasonable timeframe.
Additionally, we have released the codebase of our team to the robotics community. This open-source approach enables other teams to build upon our work, fostering collaboration and innovation in the realm of robotic soccer.
@@ Acknowledgments
Appendix A
This appendix provides additional information about team rankings, configuration parameters for the Step Baseline and skill transitions, and all the parameters and training conditions required to replicate the Sprint-Kick and the skills within the Locomotion Set. Finally, it presents additional results for the Sprint-Kick and a compilation of videos referenced throughout the paper.
Table 5: Team rankings in the RoboCup 3D Soccer Simulation League competition across the years 2008–2012. Teams sharing the same ranking are shown without a separating line.
Rank 2008 2009 2010 2011 2012
1 SEU-RedSun SEU-RedSun Apollo3D UT Austin Villa UT Austin Villa
2 WrightEagle Bold Hearts Nao Team Humboldt CIT 3D RoboCanes
3 BATS AmoiensisNQ HfutEngine3D Apollo3D Bold Hearts
4 CZU 3D HfutEngine3D Bold Hearts KylinSky3D magmaOffenburg
5 Apollo3D FUT-K SEU-RedSun Bold Hearts Apollo3D
6 Borregos3D opuCI 3D FC Portugal FC Portugal ODENS
7 Fantasia JMU 3D RoboCanes RoboCanes CIT 3D
8 MRL FC Portugal DreamWing3D SEU-RedSun FC Portugal
9 AmoiensisNQ Borregos3D opuCI 3D OxBlue3D Karachi Koalas
10 Parsian magmaOffenburg magmaOffenburg beeStanbul beeStanbul
11 NAITO-Horizon BahiaRT FUT-K magmaOffenburg FUT-K
12 OxBlue3D NomoFC beeStanbul DreamWing3D SOCIO 3D
13 FC Portugal MRL BATS Nexus 3D L3M-SIM
14 HfutEngine3D Scorpius UT Austin Villa HfutEngine3D
15 NomoFC RoboPUB Nexus 3D FUT-K
16 opuhana BATS Parsian Karachi Koalas
17 JMU 3D BMP BahiaRT Nao Team Humboldt
18 Scorpius Parsian NomoFC Rail
19 PAT Persia RoboPUB L3M-SIM
20 Strive3D Nexus 3D Alzahra Farzanegan
21 UT Austin Villa NAITO-StrikerS NomoFC
22 BahiaRT
Table 6: Team rankings in the RoboCup 3D Soccer Simulation League competition across the years 2013–2017. Teams sharing the same ranking are shown without a separating line.
Rank 2013 2014 2015 2016 2017
1 Apollo3D UT Austin Villa UT Austin Villa UT Austin Villa UT Austin Villa
2 UT Austin Villa RoboCanes FUT-K FUT-K magmaOffenburg
3 FC Portugal magmaOffenburg FC Portugal FC Portugal FUT-K
4 SEU Jolly FC Portugal BahiaRT BahiaRT RoboCanes
5 Karachi Koalas FUT-K Apollo3D magmaOffenburg AIUT3D
6 RoboCanes BahiaRT magmaOffenburg ITAndroids BahiaRT
7 magmaOffenburg SEU Jolly RoboCanes Miracle3D FC Portugal
8 Bold Hearts HfutEngine3D Nexus 3D KgpKubs KgpKubs
9 PHOTON Karachi Koalas CIT 3D HfutEngine3D ITAndroids
10 HfutEngine3D L3M-SIM HfutEngine3D Miracle3D
11 ITAndroids ODENS ITAndroids Nexus 3D
12 Paydar 3D Mithras3D Miracle3D HfutEngine3D
13 FUT-K RIC-AASTMT
14 Mithras3D
15 ODENS
16 L3M-SIM
17 BahiaRT
Table 7: Team rankings in the RoboCup 3D Soccer Simulation League competition across the years 2018–2023
Rank 2018 2019 2021 2022 2023
1 UT Austin Villa UT Austin Villa UT Austin Villa FC Portugal FC Portugal
2 magmaOffenburg magmaOffenburg magmaOffenburg magmaOffenburg magmaOffenburg
3 FC Portugal WrightOcean Apollo3D UT Austin Villa UT Austin Villa
4 BahiaRT BahiaRT WrightOcean Apollo3D ITAndroids
5 KgpKubs HfutEngine3D HfutEngine3D HfutEngine3D BahiaRT
6 Miracle3D FC Portugal FC Portugal Miracle3D WITS FC
7 ITAndroids ITAndroids Miracle3D ITAndroids
8 ITAndroids KgpKubs
9 BahiaRT WITS FC
10 KgpKubs BahiaRT
11 WITS FC
12 MIRG
A.1 Step Baseline and Transitions
@@ A.2 Sprint-Kick
step 1 Step counter [^1] In Sprint Forward and Sprint & Rotate, the counter has an upper limit of 93. A limit prevents the neural network from encountering higher values during testing than during training. The specific value was optimized as a hyperparameter, as it influences the final behavior. [3] 0 1
count 1 Counter since last side switch 0 1
z 1 Head height prediction [3] 0 1
gyro 3 Gyroscope (x-forward, y-left, z-up) (0,1,2) (-1,1,-1)
dgyro 3 $\frac{d}{dt}\begin{bmatrix}\text{gyro}_{x}&\text{gyro}_{y}&\text{gyro}_{z}\end{bmatrix}$ for $\Delta t=0.02$ s (0,1,2) (-1,1,-1)
acc 3 Accelerometer (x-forward, y-left, z-up) (0,1,2) (1,-1,1)
dacc 3 $\frac{d}{dt}\begin{bmatrix}\text{acc}_{x}&\text{acc}_{y}&\text{acc}_{z}\end{bmatrix}$ for $\Delta t=0.02$ s (0,1,2) (1,-1,1)
feet 12 (l)eft and (r)ight feet sensors with (p)oint of origin and (f)orce vector: (lpx,lpy,lpz,lfx,lfy,lfz,rpx,rpy,rpz,rfx,rfy,rfz) (6,7,8,9,10,11, 0,1,2,3,4,5) (1,-$,1$,-1,1,1,-1,1)
dfeet 12 $\frac{d}{dt}\begin{bmatrix}\text{feet}_{0}&\ldots&\text{feet}_{11}\end{bmatrix}$ for $\Delta t=0.02$ s (6,7,8,9,10,11, 0,1,2,3,4,5) (1,-$,1$,-1,1,1,-1,1)
joints 20 All joint positions excluding head and toes Sagittal reflection [^2] The symmetry operation associated with this sagittal reflection depends on the joints’ order Sagittal reflection [^2] The symmetry operation associated with this sagittal reflection depends on the joints’ order
djoints 20 $\frac{d}{dt}\begin{bmatrix}\text{joints}_{0}&\ldots&\text{joints}_{19}\end{bmatrix}$ for $\Delta t=0.02$ s Sagittal reflection [^2] The symmetry operation associated with this sagittal reflection depends on the joints’ order Sagittal reflection [^2] The symmetry operation associated with this sagittal reflection depends on the joints’ order
target [^3] The target is only included in Sprint & Rotate 1 Target direction relative to the robot’s torso 0 -1
ball [^4] The ball position is only included in Kick 3 Ball position relative to the robot’s torso     (x-forward, y-left, z-up) (0,1,2) (1,-1,1)
Table 11: Sprint-Kick action space modification sequence:
Sprint Forward, Sprint & Rotate, Kick
@@
1. Action Raw target positions for 22 joints (includes toes, excludes head joints)
2. Symmetry The positions are reflected with respect to the robot’s sagittal plane when the left leg is positioned as the front leg (see Section 6.1)
3. IIR Filter An exponential moving average filter is applied with a smoothing factor of 0.5
4. Map The raw values are mapped to the position ranges of each individual joint
5a. Bias The initial joint positions are added as a permanent bias (only for Sprint Forward)
5b. Primitive Add cycle capture primitive (only for Sprint & Rotate and Kick)
6. Conversion First, predict current joint positions by combining the last action with previous positions (the server has a 1-step delay). Then, determine joint speeds through proportional control.